# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v3

@@ 1 Introduction
One of the simplest setups of an excited state that has a clear holographic dual is the local operator quench [^1] In this paper, we will sometimes refer to the local operator quench as ‘pure-state local quench’ or simply ‘pure-state excitation’, and these names will be used interchangeably., which was first introduced in Nozaki:2014hna ; Nozaki:2014uaa with its gravitational dual given in Nozaki:2013wia in terms of a falling heavy particle in anti-de Sitter (AdS) space. CFT calculations reveal that the time evolution of entanglement entropy $\Delta S_{A}$ is sensitive to the chaotic nature of the CFT that we consider. In two-dimensional integrable CFTs, entanglement entropy increases by only a finite amount, given by the quantum dimension of the operator that was inserted to excite the system He:2014mwa ; Nozaki:2014hna ; Nozaki:2014uaa (see also Caputa:2014eta ; Guo:2015uwa ; Caputa:2015tua ; Chen:2015usa ; Nozaki:2015mca ; David:2016pzn ; Caputa:2016yzn ). In the symmetric orbifold of a compactified scalar with irrational radius, it grows like $\sim\log\log t$ Caputa:2017tju . On the other hand, for two-dimensional holographic CFTs Fitzpatrick:2014vua ; Hartman:2014oaa , which are expected to be maximally chaotic, it grows logarithmically $\sim\frac{c}{6}\log t$ when the subsystem is very large Asplund:2014coa (see also Caputa:2014vaa ; Kusuki:2017upd ; Kusuki:2018wpa ). This behavior can be explained holographically by taking into account the backreactions of heavy particles in its gravitational dual Nozaki:2013wia ; Caputa:2014vaa . Thus, we can recast the local operator quench as a problem to find the gravitational backreaction of a falling object. We can also consider the mixed-state local quench, which is a different class of local quench produced by a localized mixed state: a canonical distribution of all possible local primary operators and their descendants in a CFT Bhattacharyya:2019ifi . The holographic dual of the mixed-state local quench is given as a localized black hole in the AdS geometry. Interestingly, the mixed-state local quench is locally equivalent to the local operator quench under a certain identification of their parameters, which was pointed out in Bhattacharyya:2019ifi and will be examined more in the present paper. Moreover, a class of excitations which sit in the middle of local and global quenches has been worked out recently Goto:2021sqx ; Goto:2023wai ; Goto:2023yxb ; Kudler-Flam:2023ahk ; Mao:2024cnm ; Miyata:2024gvr ; Bai:2024azk .
@@ 2.1 HHLL approximation of conformal blocks
where $H$ is the CFT Hamiltonian and $\delta$ is the regularization parameter. Taking subsystem $A$ to be a single interval $[a,b]$, we are interested in computing $\Delta S_{A}$, the entanglement entropy of this excited state relative to the vacuum. Morally speaking, the operator insertion produces an entangled pair of excitations that travel in opposite directions at the speed of light. The location of the pair in relation to the subsystem is what determines $\Delta S_{A}$. While the result for this particular setup is already known, the purpose of this section is to review the method that is employed in the computation and to apply it on a setup that involves an additional excitation in the form of a mixed-state local operator insertion.
@@ 2.2 Computation of entanglement entropy
$\theta$ $=2\pi n+2\delta\left(\frac{1}{a-t}-\frac{1}{b-t}\right)+O(\delta^{2}),$ (2.18)
$\bar{\theta}$ $=2\pi\bar{n}+2\delta\left(\frac{1}{a+t}-\frac{1}{b+t}\right)+O(\delta^{2}),$ (2.19)
where $|n|,|\bar{n}|\leq 1$. The correct choice of branch will need to be determined on physical grounds, as both 1z1-z and $1-\bar{z}$ exhibit monodromic behavior depending on the position of the operator insertion relative to the subsystem.
@@
Let us first consider the case where the operator insertion lies outside the subsystem, $ab>0$. There are two possible configurations: $a,b>0$ (which we shall call ‘case I’) and $a,b<0$ (‘case III’). Without loss of generality, assume case I. For $t<a$, causality tells us that we should expect $\Delta S_{A}=0$, which can only be achieved if $n=\bar{n}=0$. Indeed, starting from this branch and letting $z$, $\bar{z}$ evolve in time, we see that for $a<t<b$, $n=1$ while $\bar{n}=0$, which gives us finite $\Delta S_{A}$
@@
For $t>b$, we are once again back at $n=\bar{n}=0$, giving us $\Delta S_{A}=0$ as expected.
@@
Let us now consider the case where the operator insertion lies inside the subsystem, $a<0<b$ (‘case II’). Once again, without loss of generality, assume $|a|<b$. Causality mandates that $\Delta S_{A}=0$ when $t<|a|$ and $t>b$, which implies $n=\bar{n}=0$ for those times. However, there seemingly arises a problem of consistency when we impose those conditions. If we impose $n=\bar{n}=0$ for $t>b$ and let $z,\bar{z}$ evolve backwards in time, we see that for $|a|<t<b$, $n=1,\bar{n}=0$, and for $t<|a|$, $n=\bar{n}=1$. In order to impose $n=\bar{n}=0$ for $t<|a|$, we have to consider another channel by replacing $1-z\to e^{-2\pi i}(1-z)$ and $1-\bar{z}\to e^{2\pi i}(1-\bar{z})$, which corresponds to a shift of $(n,\bar{n})\mapsto(n-1,\bar{n}-1)$. At any time, the correct channel should always minimize the entanglement entropy, so
@@
where $i$ labels the channels. Applying this prescription, we get
@@
$\Delta S_{A}=\frac{c}{6}\log\left(\frac{\sin\left(\alpha_{H}\pi\right)t}{\alpha_{H}\delta}\right)$ (2.23)
For later convenience, let us name the channel that imposes $n=\bar{n}=0$ at late times ‘channel 1’ and the other channel ‘channel 2’. By comparing the functional form of (2.20) and (2.22), we can see that channel 1 is the only dominant channel and channel 2 is irrelevant when the insertion is outside the subsystem.
@@
    • the left endpoint $a$ of the subsystem: $(n,\bar{n})\mapsto(n+1,\bar{n})$,
    • the right endpoint $b$ of the subsystem: $(n,\bar{n})\mapsto(n-1,\bar{n})$.
@@
    • the right endpoint $b$ of the subsystem: $(n,\bar{n})\mapsto(n,\bar{n}+1)$,
    • the left endpoint $a$ of the subsystem: $(n,\bar{n})\mapsto(n,\bar{n}-1)$.
The different channels are simply there to set different initial conditions for $(n,\bar{n})$. For case I, we only see monodromic behavior from 1z1-z, whereas for case III, we only see monodromic behavior from $1-\bar{z}$. In either of these cases, $(n,\bar{n})$ at $t=0$ and at large $t$ are always the same, which is why there is only one dominant channel. For case II, we see monodromic behavior from both 1z1-z and $1-\bar{z}$, but they revolve once and do not revolve back to $1-z=1-\bar{z}=1$. So $(n,\bar{n})$ at $t=0$ and at large $t$ are always different, which necessitates contribution from a secondary channel.
Another related observation that we can make from (2.15) is that only the variables that exhibit monodromy matter. Provided that at $n=0$ or $\bar{n}=0$, $\theta,\bar{\theta}$ can be treated as small angles, we see that $\Delta S_{A}$ assumes a nonzero value only when $n\neq 0$ or $\bar{n}\neq 0$. So in case I (III), we should only care about the behavior of $z$ ($\bar{z}$); we do not have to carry out explicit calculations to find out the value $\bar{\theta}$ ($\theta$) assumes. Obviously, from our previous arguments, case II is a lot more subtle and both $z$ and $\bar{z}$ have to be carefully tracked. We shall see in section 3 that in the double excitation setup, the assumption of $\theta,\bar{\theta}$ being small at $n=0$ or $\bar{n}=0$ starts to fail in certain case II configurations, especially when the regulators for the two excitations become comparable to each other in size.
@@ 2.3 Holographic interpretation
$\theta=\Delta\tilde{\phi}-\Delta\tilde{t}$,
$\bar{\theta}=\Delta\tilde{\phi}+\Delta\tilde{t}$, (2.28)
then these agree with the angles in (2.14). Note that $\Delta\tilde{\phi}$ is meaningful only in the range $0<\Delta\tilde{\phi}<2\pi$ as $\Delta\tilde{\phi}=0$ and $\Delta\tilde{\phi}=2\pi$ are equivalent due to the periodicity of the deficit angle. The entanglement entropy comes out as
@@
$=\frac{1}{4G_{N}}\min\left[D_{12}(\theta,\bar{\theta}),\ D_{12}(2\pi-\bar{\theta},2\pi-\theta)\right].$ (2.29)
Indeed, this reproduces the result (2.15) obtained from the HHLL calculation of holographic CFTs.
@@ 3.3 Insertion of local operator excitation
In the previous subsection, we found that mixed-state excitations correspond to introducing an identified set of finite-sized holes in the Euclidean geometry of the theory. Another important takeaway from the section was the special property of holographic theories that allows us to treat this geometry, which naively is conformally equivalent to a torus, as instead being equivalent to an infinitely long cylinder. We can effectively decompactify one of the directions of spacetime, giving rise to two distinct phases — BTZ black hole and thermal AdS — depending on our choice of direction. What we would like to do now is to consider local operator excitations, as done in previous sections, on this geometry. On a technical level this calculation is similar to the one carried out in Caputa:2014eta , where the local operator quench was studied at finite temperature. We shall mainly be interested in the BTZ phase, where the spatial direction is decompactified, as it gives physically nontrivial results. The thermal AdS phase is discussed in appendix A, where we shall see that results are similar to what we have already seen for single local operator excitations.
@@
Figure 4: Our setup of interest depicted on the Lorentzian plane. The local operator excitation by $\cal O$} produces the entangled pair of modes which propagate in the left and right direction at the speed of light. When $t_{M}+x_{P}=0$, they collides directly.
@@
In addition to the mixed-state excitation at $(x,t)=(0,t_{M})$, consider adding a local operator excitation $\mathcal{O}$ at $(x,t)=(x_{P},0)$, as depicted in figure 4. Without loss of generality, assume $x_{P}\leq 0$ throughout the paper. When this is passed on to the Euclidean plane, the locations of pure-state and mixed-state excitations are as shown in figure 2. The local operator excitation by $\cal O$} produces a entangled pair of modes which propagate in the left and right direction at the speed of light. When $t_{M}+x_{P}=0$, they precisely collides with each other. The main interest in this paper is how the entanglement entropy is affected by this collision as we will analyze below.
@@
Calculation of entanglement entropy requires the computation of four-point function $\langle\mathcal{O}^{\dagger}\mathcal{O}\sigma_{n}\bar{\sigma}_{n}\rangle$, which can be done via the HHLL approximation as previously seen in section 2.1. Our strategy is therefore straightforward: we must first conformally transform the two-hole Euclidean plane into an annulus where the time direction is compactified and the space direction extends radially. This was done partially in the previous subsection, where we made use of the conformal transformation that mapped the two-hole geometry onto a torus, or more accurately a finite strip with opposite sides identified. A further application of the exponential map with the correct power should yield the desired annulus. With the space direction extending radially, we can now reap the benefits of working with a holographic theory and ignore the two boundaries of the annulus, which means that the geometry is now the full Euclidean plane. At this point, we can simply apply the method of HHLL approximation and compute the four-point function by tracking the behavior of cross ratios of insertion points.
@@
Notice that we have two different UV cutoffs. The UV cutoffs of the local operator and mixed-state excitation are $\delta$ and $s$, respectively. Depending on how we take the limits $\delta,s\to 0$, factors such as $\zeta_{\mathcal{O}}$ and $\zeta_{\mathcal{O}^{\dagger}}$ will assume different values. We shall treat both cases of $\delta<s$ and $s<\delta$, the former being more straightforward than the latter. The apparent issue with the $s<\delta$ regime is that $\zeta_{\mathcal{O}},\zeta_{\mathcal{O}^{\dagger}}\to(-1)^{\frac{2\pi i}{\beta}}$ as $s\ll\delta$, but there is no way of deciding their logarithmic branches. We will return to this issue in section 4.3, where we discuss this ambiguity in detail and see how different choices affect the computation of $\Delta S_{A}$.

### v4

@@ 1 Introduction
One of the simplest setups of an excited state that has a clear holographic dual is the local operator quench [^1] In this paper, we will sometimes refer to the local operator quench as ‘pure-state local quench’ or simply ‘pure-state excitation’, and these names will be used interchangeably., which was first introduced in Nozaki:2014hna; Nozaki:2014uaa with its gravitational dual given in Nozaki:2013wia in terms of a falling heavy particle in anti-de Sitter (AdS) space. CFT calculations reveal that the time evolution of entanglement entropy $\Delta S_{A}$ is sensitive to the chaotic nature of the CFT that we consider. In two-dimensional integrable CFTs, entanglement entropy increases by only a finite amount, given by the quantum dimension of the operator that was inserted to excite the system He:2014mwa; Nozaki:2014hna; Nozaki:2014uaa (see also Caputa:2014eta; Guo:2015uwa; Caputa:2015tua; Chen:2015usa; Nozaki:2015mca; David:2016pzn; Caputa:2016yzn). In the symmetric orbifold of a compactified scalar with irrational radius, it grows like $\sim\log\log t$ Caputa:2017tju. On the other hand, for two-dimensional holographic CFTs Fitzpatrick:2014vua; Hartman:2014oaa, which are expected to be maximally chaotic, it grows logarithmically $\sim\frac{c}{6}\log t$ when the subsystem is very large Asplund:2014coa (see also Caputa:2014vaa; Kusuki:2017upd; Kusuki:2018wpa). This behavior can be explained holographically by taking into account the backreactions of heavy particles in its gravitational dual Nozaki:2013wia; Caputa:2014vaa. Local operator quenches in higher-dimensional CFTs have also recently been studied in Bianchi:2025fzs. Another setup we can consider is the mixed-state local quench, which is a different class of local quench produced by a localized mixed state: a canonical distribution of all possible local primary operators and their descendants in a CFT Bhattacharyya:2019ifi. The holographic dual of the mixed-state local quench is given as a localized black hole in the AdS geometry. Interestingly, the mixed-state local quench is locally equivalent to the local operator quench under a certain identification of their parameters, which was pointed out in Bhattacharyya:2019ifi and will be examined more in the present paper. Moreover, a class of excitations which sit in the middle of local and global quenches has been worked out recently Goto:2021sqx; Goto:2023wai; Goto:2023yxb; Kudler-Flam:2023ahk; Mao:2024cnm; Miyata:2024gvr; Bai:2024azk.
@@ 2.1 HHLL approximation of conformal blocks
where $H$ is the CFT Hamiltonian and $\delta$ is the regularization parameter. Taking subsystem $A$ to be a single interval $[a,b]$, we are interested in computing $\Delta S_{A}$, the entanglement entropy of this excited state relative to the vacuum. Morally speaking, the operator insertion produces an entangled pair of excitations that travel in opposite directions at the speed of light. The location of the pair in relation to the subsystem is what determines $\Delta S_{A}$. Specifically, when both modes of the pair are inside/outside the subsystem, entanglement entropy is expected to vanish. On the other hand, when only one of the modes is inside the subsystem, there is entanglement between the subsystem and its complement, giving rise to nontrivial entanglement entropy $\Delta S_{A}\neq 0$. This physical interpretation of the local quench is often referred to as the quasiparticle picture. While the analytical result for this particular setup is already known, the purpose of this section is to review the method that is employed in the computation and to apply it on a setup that involves an additional excitation in the form of a mixed-state local operator insertion.
@@ 2.2 Computation of entanglement entropy
$\theta$ $=2\pi m+2\delta\left(\frac{1}{a-t}-\frac{1}{b-t}\right)+O(\delta^{2}),$ (2.18)
$\bar{\theta}$ $=2\pi\bar{m}+2\delta\left(\frac{1}{a+t}-\frac{1}{b+t}\right)+O(\delta^{2}),$ (2.19)
where $|m|,|\bar{m}|\leq 1$ (as $|\theta|,|\bar{\theta}|<2\pi$). The correct choice of branch will need to be determined on physical grounds, as both 1z1-z and $1-\bar{z}$ exhibit monodromic behavior depending on the position of the operator insertion relative to the subsystem.
@@
Let us first consider the case where the operator insertion lies outside the subsystem, $ab>0$. There are two possible configurations: $a,b>0$ (which we shall call ‘case I’) and $a,b<0$ (‘case III’). Without loss of generality, assume case I. For $t<a$, causality tells us that we should expect $\Delta S_{A}=0$, which can only be achieved if $m=\bar{m}=0$. Indeed, starting from this branch and letting $z$, $\bar{z}$ evolve in time, we see that for $a<t<b$, $m=1$ while $\bar{m}=0$, which gives us finite $\Delta S_{A}$:
@@
For $t>b$, we are once again back at $m=\bar{m}=0$, giving us $\Delta S_{A}=0$ as expected. In terms of the quasiparticle picture, while the left-moving mode is always outside the subsystem, the right-moving mode enters the subsystem at $t=a$ and leaves at $t=b$, matching the expectation that $\Delta S_{A}$ is non-vanishing for $a<t<b$.
@@
Let us now consider the case where the operator insertion lies inside the subsystem, $a<0<b$ (‘case II’). Once again, without loss of generality, assume $|a|<b$. The quasiparticle picture tells us that $\Delta S_{A}=0$ when $t<|a|$ (as both modes are inside the subsystem) and $t>b$ (as both modes are outside the subsystem), which implies $m=\bar{m}=0$ for those times. However, there seemingly arises a problem of consistency when we impose those conditions. If we impose $m=\bar{m}=0$ for $t>b$ and let $z,\bar{z}$ evolve backwards in time, we see that for $|a|<t<b$, $m=1,\bar{m}=0$, and for $t<|a|$, $m=\bar{m}=1$. This gives us
$\Delta S_{A,1}=\begin{cases}\frac{c}{6}\log\left(\frac{\sin^{2}\left(\alpha_{H}\pi\right)(t^{2}-a^{2})(b^{2}-t^{2})}{\alpha_{H}^{2}\delta^{2}(b-a)^{2}}\right),&t<|a|\\ \frac{c}{6}\log\left(\frac{\sin\left(\alpha_{H}\pi\right)(t-a)(b-t)}{\alpha_{H}\delta(b-a)}\right),&|a|<t<b\\ 0,&t>b\end{cases}$ (2.21)
where 1 in the subscript of $\Delta S_{A,1}$ implies we are considering a particular channel that fixes $(m,\bar{m})$ at some point in time. In order to impose $m=\bar{m}=0$ for $t<|a|$, we have to consider another channel by replacing $1-z\mapsto e^{-2\pi i}(1-z)$ and $1-\bar{z}\mapsto e^{2\pi i}(1-\bar{z})$, which corresponds to a shift of $(m,\bar{m})\mapsto(m-1,\bar{m}-1)$. The second channel gives
$\Delta S_{A,2}=\begin{cases}0,&t<|a|\\ \frac{c}{6}\log\left(\frac{\sin\left(\alpha_{H}\pi\right)(t+a)(b+t)}{\alpha_{H}\delta(b-a)}\right),&|a|<t<b\\ \frac{c}{6}\log\left(\frac{\sin^{2}\left(\alpha_{H}\pi\right)(t^{2}-a^{2})(b^{2}-t^{2})}{\alpha_{H}^{2}\delta^{2}(b-a)^{2}}\right).&t>b\end{cases}$ (2.22)
Among these two candidate channels, the correct choice at any given time should always minimize the entanglement entropy, so
@@
where $i$ labels the channel. Applying this prescription, we finally obtain
@@
$\Delta S_{A}=\frac{c}{6}\log\left(\frac{\sin\left(\alpha_{H}\pi\right)t}{\alpha_{H}\delta}\right).$ (2.25)
For later convenience, let us call the channel that imposes $m=\bar{m}=0$ at late times ‘channel 1’ and the other channel ‘channel 2’. By comparing the functional form of (2.20) and (2.24), we can see that channel 1 is the only dominant channel and channel 2 is irrelevant when the insertion is outside the subsystem.
@@
    • the left endpoint $a$ of the subsystem: $(m,\bar{m})\mapsto(m+1,\bar{m})$,
    • the right endpoint $b$ of the subsystem: $(m,\bar{m})\mapsto(m-1,\bar{m})$.
@@
    • the right endpoint $b$ of the subsystem: $(m,\bar{m})\mapsto(m,\bar{m}+1)$,
    • the left endpoint $a$ of the subsystem: $(m,\bar{m})\mapsto(m,\bar{m}-1)$.
The different channels are simply there to set different initial conditions for $(m,\bar{m})$. For case I, we only see monodromic behavior from 1z1-z, whereas for case III, we only see monodromic behavior from $1-\bar{z}$. In either of these cases, $(m,\bar{m})$ at $t=0$ and at large $t$ are always the same, which is why there is only one dominant channel. For case II, we see monodromic behavior from both 1z1-z and $1-\bar{z}$, but they revolve once and do not revolve back to $1-z=1-\bar{z}=1$. So $(m,\bar{m})$ at $t=0$ and at large $t$ are always different, which necessitates contribution from a secondary channel.
Another related observation that we can make from (2.15) is that only the variables that exhibit monodromy matter. Provided that $\theta,\bar{\theta}$ can be treated as small angles at $m=0$ or $\bar{m}=0$, we see that $\Delta S_{A}$ assumes a nonzero value only when $m\neq 0$ or $\bar{m}\neq 0$. So in case I (III), we should only care about the behavior of $z$ ($\bar{z}$); we do not have to carry out explicit calculations to find out the value $\bar{\theta}$ ($\theta$) assumes. Obviously, from our previous arguments, case II is a lot more subtle and both $z$ and $\bar{z}$ have to be carefully tracked. We shall see in section 3 that in the double excitation setup, the assumption of $\theta,\bar{\theta}$ being small at $m=0$ or $\bar{m}=0$ starts to fail in certain case II configurations, especially when the regulators for the two excitations become comparable to each other in size.
@@ 2.3 Holographic interpretation
$\theta=\Delta\tilde{\phi}-\Delta\tilde{t},\qquad\bar{\theta}=\Delta\tilde{\phi}+\Delta\tilde{t}$, (2.30)
then these agree with the angles in (2.14). Note that $\Delta\tilde{\phi}$ is meaningful only in the range $0<\Delta\tilde{\phi}<2\pi$ as $\tilde{\phi}\sim\tilde{\phi}+2\pi$. The entanglement entropy comes out as
@@
$=\frac{1}{4G_{N}}\min\left[D_{12}(\theta,\bar{\theta}),\ D_{12}(2\pi-\bar{\theta},2\pi-\theta)\right]$
$=\frac{1}{4G_{N}}\min\left[D_{12}(\theta,\bar{\theta}),\ D_{12}(\theta-2\pi,\bar{\theta}-2\pi)\right],$ (2.31)
where the second geodesic is obtained by taking the subsystem that wraps around the other side of the cylinder $\Delta\tilde{\phi}\mapsto 2\pi-\Delta\tilde{\phi}$. From (2.3), it is evident that taking the second geodesic is equivalent to choosing channel 2, replacing $\theta\mapsto\theta-2\pi$, $\bar{\theta}\mapsto\bar{\theta}-2\pi$ and hence $(m,\bar{m})\mapsto(m-1,\bar{m}-1)$. It follows that $\Delta S_{A}=S_{A}(\alpha)-S_{A}(1)$ reproduces the result (2.15) obtained from the HHLL calculation of holographic CFTs.
@@ 3.3 Insertion of local operator excitation
In the previous subsection, we found that mixed-state excitations correspond to introducing an identified set of finite-sized holes in the Euclidean geometry of the theory. Another important takeaway from the section was the special property of holographic theories that allows us to treat this geometry, which naively is conformally equivalent to a torus, as instead being equivalent to an infinitely long cylinder. We can effectively decompactify one of the directions of spacetime, giving rise to two distinct phases — BTZ black hole and thermal AdS — depending on our choice of direction. What we would like to do now is to consider local operator excitations on this geometry. On a technical level, this calculation is similar to the one carried out in Caputa:2014eta, where the local operator quench was studied at finite temperature. In this paper, we shall mainly be interested in the BTZ phase, where the spatial direction is decompactified, as it gives physically nontrivial results. The thermal AdS phase is discussed in appendix A, where we shall see that results are similar to what we have already seen for single local operator excitations.
At this point, it is natural to question whether the cylindrical approximation of the geometry holds up in the presence of heavy operators. In section 4.1, we will argue that the approximation is valid when the regulators $\delta,s$ of quenches are not of the same order, i.e. $\delta/s\ll 1$ or $\delta/s\gg 1$. We shall therefore limit our discussion in this section to those two particular limits of regulators. Taking these limits also allows us to compute simple analytical expressions for entanglement entropy — something we always like.
@@
Figure 4: Our setup of interest depicted on the Lorentzian plane. The local operator excitation by $\cal O$} produces an entangled pair of modes that propagate in opposite directions at the speed of light. When $t_{M}+x_{P}=0$, one of the modes directly collides with the mixed-state excitation.
@@
In addition to the mixed-state excitation at $(x,t)=(0,t_{M})$ with $t_{M}>0$, consider adding a local operator excitation $\mathcal{O}$ at $(x,t)=(x_{P},0)$, as depicted in figure 4. When $x_{P}\pm t_{M}=0$, one of the entangled pair of modes produced by the local operator excitation collides with the mixed-state excitation, where we expect an interesting interplay. Without loss of generality, assume $x_{P}\leq 0$ hereafter (so lightlike separation is only attainable when $x_{P}+t_{M}=0$). Once we pass this on to the Euclidean plane, the locations of pure-state and mixed-state excitations are as shown in figure 2.
@@
Calculation of entanglement entropy requires the computation of four-point function $\langle\mathcal{O}^{\dagger}\mathcal{O}\sigma_{n}\bar{\sigma}_{n}\rangle$, which can be done via the HHLL approximation as previously seen in section 2.1. Our strategy is therefore straightforward: we must first conformally transform the two-hole geometry in figure 2 into an annulus where the time direction is compactified and the space direction extends radially. This was done partially in the previous subsection, where we made use of the conformal transformation that mapped the two-hole geometry onto a finite strip. A further application of the exponential map with the correct power should yield the desired annulus. With the space direction extending radially, we can now reap the benefits of working with a holographic theory and ignore the two boundaries of the annulus, which means that the geometry is now the full Euclidean plane. At this point, we can simply apply the method of HHLL approximation and compute the four-point function by tracking the behavior of cross ratios of insertion points.
@@
Notice that we have two different UV cutoffs $\delta,s$, corresponding to cutoffs for the local operator and mixed-state quench respectively. Depending on how we take the limits $\delta,s\to 0$, factors such as $\zeta_{\mathcal{O}}$ and $\zeta_{\mathcal{O}^{\dagger}}$ will assume different values. We shall treat both cases of $\delta<s$ and $s<\delta$, the former being more straightforward than the latter. The apparent issue with the $s<\delta$ regime is that $\zeta_{\mathcal{O}},\zeta_{\mathcal{O}^{\dagger}}\to(-1)^{\frac{2\pi i}{\beta}}$ as $s\ll\delta$, but there is no way of deciding their logarithmic branches. We will return to this issue in section 4.3, where we discuss this ambiguity in detail and see how different choices affect the computation of $\Delta S_{A}$.